50 ml of pure and dry oxygen gas was subjected to electric discharge and on cooling to original temperature the volume of ozonized oxygen was found to be 47 ml. The gas was then brought in contact with turpentine oil the remaining gas occupied a volume of 41 ml. Find molecular formula of ozone

Grade:11

1 Answers

Arun
3 years ago
The reaction of oxygen to ozone can be shown as:
 
3O_2 -> 2O_3
 
Now, for 90 ml of O_2 volume of Ozone = 90*2/3 = 60 ml
 
Only 10% O_3 is converted and can be shown as:
 
Volume of ozone = 10 % of 60 = 6 ml.
 
Hence,
 
AFter some time the remaining volume of ozone = 6 mL
 
Volume of oxygen = 90 – 6 = 84 mL